# On-Device Language Models: A Comprehensive Review

### v1

@@ 2.1 Evolution of On-Device LLMs
  • Nexa AI’s Octopus series (Chen & Li (2024a; b; c))
  • ModelBest’s MiniCPM series (Hu et al. (2024b); Tsinghua University (2024))
  • Google’s Gemma series (Team et al. (2024); Google (2024a))
  • Apple’s OpenELM (Mehta et al. (2024)) and DataComp-LM (Li et al. (2024a))
  • AI2’s OLMo (Soldaini et al. (2024); Groeneveld et al. (2024))
@@ 2.4 Limitations of Cloud-Based LLM Inference and Advantages of On-Device Inference
Although cloud-based LLMs offer powerful capabilities, they come with certain drawbacks, including potential latency issues (Wang et al., 2024b) and data concerns due to their dependency on networks. Hence, the concept of on-device deployment through edge computing has emerged to reduce latency and safeguard user data (Gerganov (2023)). Processing occurs locally, eliminating the need for data transmission. Moreover, the proliferation of customized hardware accelerators on mobile devices has made it feasible to run large LLMs with billions of parameters directly on devices.
@@ 3.1 Architectural Design Principles and Innovations for On-Device LLMs
MobileLLM (Liu et al. (2024c)) High accuracy, optimized for sub-billion parameter models Embedding sharing, grouped-query attention Reduced model size due to deep and thin structures
EdgeShard (Zhang et al. (2024b)) Up to 50% latency reduction, 2× throughput improvement Collaborative edge-cloud computing, optimal shard placement Distributed model components reduce individual device load
LLMCad (Xu et al. (2023)) Up to 9.3× speedup in token generation Generate-then-verify, token tree generation Smaller LLM for token generation, larger LLM for verification
Any-Precision LLM (Park et al. (2024)) Supports multiple precisions efficiently Post-training quantization, memory-efficient design Substantial memory savings with versatile model precisions
Breakthrough Memory (Kim et al. (2024c)) Up to 4.5× performance improvement PIM and PNM technologies enhance memory processing Enhanced memory bandwidth and capacity
MELTing Point (Laskaridis et al. (2024)) Provides systematic performance evaluation Analyzes impacts of quantization, efficient model evaluation Evaluates memory and computational efficiency trade-offs
LLMaaS on MD (Yin et al. (2024)) Reduces context switching latency significantly Stateful execution, fine-grained KV cache compression Efficient memory management with tolerance-aware compression and swapping
LocMoE (Li et al. (2024b)) Reduces training time per epoch by up to 22.24% Orthogonal gating weights, locality-based expert regularization Minimizes communication overhead with group-wise All-to-All and recompute pipeline
EdgeMoE (Yi et al. (2023)) Significant performance improvements on edge devices Expert-wise bitwidth adaptation, preloading experts Efficient memory management through expert-by-expert computation reordering
JetMoE (Shen et al. (2024)) Outperforms Llama2-7B and 13B-Chat with fewer parameters Reduces inference computation by  70% using sparse activation 8B total parameters, only 2B activated per input token
@@ 5.1 Popular On-Device LLMs Framework
      • Description: Llama.cpp (Gerganov (2023)) is a C/C++ library designed for efficient inference of large language models on a broad range of hardware platforms. It supports integer quantization, GPU acceleration, and CPU+GPU hybrid inference.
@@
      • Description: MNN (Alibaba (2024)) leverages Mobile Neural Network technology for efficient LLM inference on various platforms, optimized for mobile devices with dynamic inputs and multimodal interactions.
@@
      • Description: PowerInfer (Song et al. (2023)) and PowerInfer2 (Xue et al. (2024b)) is a high-speed inference engine optimized for deploying LLMs on PCs with consumer-grade GPUs, utilizing a locality-centric design.
@@
      • Description: ExecuTorch (PyTorch (2024)) is part of the PyTorch Edge ecosystem, designed for deploying PyTorch models efficiently on edge devices like mobile phones and wearables.
@@
      • Description: Developed by Google, MediaPipe (AI (2024b)) is a framework for building and deploying multimodal machine learning pipelines involving video, audio, and other time-series data.
@@
      • Description: MLC-LLM (team (2023)) is a machine learning compiler and high-performance deployment engine, supporting universal LLM deployment on edge devices and in cloud environments.
@@
      • Description: VLLM (Team (2024)) is optimized for edge-cloud environments, supporting advanced quantization methods for efficient key and value memory management during inference.
@@
      • Description: OpenLLM (BentoML (2024)) enables the deployment of various open-source LLMs as OpenAI-compatible API endpoints, optimized for high throughput and streamlined cloud deployment.
@@ 6.2 Applications of On-Device LLMs
    Similar capabilities can also be used for sign language recognition, and there are projects that use the ChatGPT model for sign language translation (Sincan et al. (2024)). In comparison, the on-device model can generate text translations corresponding to sign language with lower latency and ensure its offline availability.
@@
    Autonomous Vehicles: Using language models to drive autonomous cars may be an ideal future, but we already have examples of this today. DriveVLM Dual is a system that combines autonomous driving technology with a large-scale visual language model (VLM) to improve the understanding of complex and long-tail scenes in urban environments. The system uses language to describe the driving environment and identify key objects in the scene. It gradually develops a plan from meta-action and decision descriptions to waypoints. DriveVLM surpasses existing state-of-the-art methods on both public benchmarks and the researchers’ own benchmarks, especially in handling complex and dynamic scenes. Excitingly, DriveVLM can be deployed locally on the car, which also provides convenience for its immediate response (Tian et al. (2024)).
@@ 7 Future Directions and Open Challenges
As on-device LLMs continue to evolve, several vital areas emerge as promising future research and development directions. The field of on-device LLMs is rapidly advancing, driven by the increasing demand for 1) data security, 2) low-latency, and 3) personalized AI experiences on edge devices. This progress is exemplified by recent developments such as TinyLlama (Zhang et al. (2024c)), MobileVLM (Murthy et al. (2024); Chu et al. (2024)), and novel approaches like the OpenELM (Mehta et al. (2024)). However, deploying LLMs on resource-constrained devices presents unique challenges that differ significantly from traditional cloud-based implementations. These challenges span multiple areas, including model compression, efficient inference, security, energy efficiency, and seamless integration with diverse hardware platforms. Moreover, the dynamic nature of edge environments and the need for continuous adaptation introduce additional complexities that must be considered. We outline the most pressing challenges and opportunities in advancing the field of LLMs on-device. By identifying these key areas and stimulating innovation in developing more capable, efficient, and reliable on-device language models, we aim to provide insights for future research efforts. We should notice that the challenges and opportunities discussed here are interconnected: the progress in one area often has implications for others. Therefore, a holistic approach that considers the interplay between different aspects of on-device LLM deployment is essential for achieving significant advancements in the field. We delve into the current state of research, identifying key challenges and proposing potential directions for future work, summarized in Fig. 6. By addressing these challenges, researchers and practitioners can push the boundaries of what is possible with on-device LLMs, ultimately leading to more intelligent, efficient, and user-centric computing experiences across various applications and domains.

### v2

@@ 2.1 Evolution of On-Device LLMs
  • Nexa AI’s Octopus series (Chen & Li, 2024a; b; c)
  • ModelBest’s MiniCPM series (Hu et al., 2024b; Tsinghua University, 2024)
  • Google’s Gemma series (Team et al., 2024; Google, 2024a)
  • Apple’s OpenELM (Mehta et al., 2024) and DataComp-LM (Li et al., 2024a)
  • AI2’s OLMo (Soldaini et al., 2024; Groeneveld et al., 2024)
@@ 2.4 Limitations of Cloud-Based LLM Inference and Advantages of On-Device Inference
Although cloud-based LLMs offer powerful capabilities, they come with certain drawbacks, including potential latency issues (Wang et al., 2024b) and data concerns due to their dependency on networks. Hence, the concept of on-device deployment through edge computing has emerged to reduce latency and safeguard user data (Gerganov, 2023). Processing occurs locally, eliminating the need for data transmission. Moreover, the proliferation of customized hardware accelerators on mobile devices has made it feasible to run large LLMs with billions of parameters directly on devices.
@@ 3.1 Architectural Design Principles and Innovations for On-Device LLMs
MobileLLM (Liu et al., 2024c) High accuracy, optimized for sub-billion parameter models Embedding sharing, grouped-query attention Reduced model size due to deep and thin structures
EdgeShard (Zhang et al., 2024b) Up to 50% latency reduction, 2× throughput improvement Collaborative edge-cloud computing, optimal shard placement Distributed model components reduce individual device load
LLMCad (Xu et al., 2023) Up to 9.3× speedup in token generation Generate-then-verify, token tree generation Smaller LLM for token generation, larger LLM for verification
Any-Precision LLM (Park et al., 2024) Supports multiple precisions efficiently Post-training quantization, memory-efficient design Substantial memory savings with versatile model precisions
Breakthrough Memory (Kim et al., 2024c) Up to 4.5× performance improvement PIM and PNM technologies enhance memory processing Enhanced memory bandwidth and capacity
MELTing Point (Laskaridis et al., 2024) Provides systematic performance evaluation Analyzes impacts of quantization, efficient model evaluation Evaluates memory and computational efficiency trade-offs
LLMaaS on MD (Yin et al., 2024) Reduces context switching latency significantly Stateful execution, fine-grained KV cache compression Efficient memory management with tolerance-aware compression and swapping
LocMoE (Li et al., 2024b) Reduces training time per epoch by up to 22.24% Orthogonal gating weights, locality-based expert regularization Minimizes communication overhead with group-wise All-to-All and recompute pipeline
EdgeMoE (Yi et al., 2023) Significant performance improvements on edge devices Expert-wise bitwidth adaptation, preloading experts Efficient memory management through expert-by-expert computation reordering
JetMoE (Shen et al., 2024) Outperforms Llama2-7B and 13B-Chat with fewer parameters Reduces inference computation by  70% using sparse activation 8B total parameters, only 2B activated per input token
@@ 5.1 Popular On-Device LLMs Framework
      • Description: Llama.cpp (Gerganov, 2023) is a C/C++ library designed for efficient inference of large language models on a broad range of hardware platforms. It supports integer quantization, GPU acceleration, and CPU+GPU hybrid inference.
@@
      • Description: MNN (Alibaba, 2024) leverages Mobile Neural Network technology for efficient LLM inference on various platforms, optimized for mobile devices with dynamic inputs and multimodal interactions.
@@
      • Description: PowerInfer (Song et al., 2023) and PowerInfer2 (Xue et al., 2024b) is a high-speed inference engine optimized for deploying LLMs on PCs with consumer-grade GPUs, utilizing a locality-centric design.
@@
      • Description: ExecuTorch (PyTorch, 2024) is part of the PyTorch Edge ecosystem, designed for deploying PyTorch models efficiently on edge devices like mobile phones and wearables.
@@
      • Description: Developed by Google, MediaPipe (AI, 2024b) is a framework for building and deploying multimodal machine learning pipelines involving video, audio, and other time-series data.
@@
      • Description: MLC-LLM (team, 2023) is a machine learning compiler and high-performance deployment engine, supporting universal LLM deployment on edge devices and in cloud environments.
@@
      • Description: VLLM (Team, 2024) is optimized for edge-cloud environments, supporting advanced quantization methods for efficient key and value memory management during inference.
@@
      • Description: OpenLLM (BentoML, 2024) enables the deployment of various open-source LLMs as OpenAI-compatible API endpoints, optimized for high throughput and streamlined cloud deployment.
@@ 6.2 Applications of On-Device LLMs
    Similar capabilities can also be used for sign language recognition, and there are projects that use the ChatGPT model for sign language translation (Sincan et al., 2024). In comparison, the on-device model can generate text translations corresponding to sign language with lower latency and ensure its offline availability.
@@
    Autonomous Vehicles: Using language models to drive autonomous cars may be an ideal future, but we already have examples of this today. DriveVLM Dual is a system that combines autonomous driving technology with a large-scale visual language model (VLM) to improve the understanding of complex and long-tail scenes in urban environments. The system uses language to describe the driving environment and identify key objects in the scene. It gradually develops a plan from meta-action and decision descriptions to waypoints. DriveVLM surpasses existing state-of-the-art methods on both public benchmarks and the researchers’ own benchmarks, especially in handling complex and dynamic scenes. Excitingly, DriveVLM can be deployed locally on the car, which also provides convenience for its immediate response (Tian et al., 2024).
@@ 7 Future Directions and Open Challenges
As on-device LLMs continue to evolve, several vital areas emerge as promising future research and development directions. The field of on-device LLMs is rapidly advancing, driven by the increasing demand for 1) data security, 2) low-latency, and 3) personalized AI experiences on edge devices. This progress is exemplified by recent developments such as TinyLlama (Zhang et al., 2024c), MobileVLM (Murthy et al., 2024; Chu et al., 2024), and novel approaches like the OpenELM (Mehta et al., 2024). However, deploying LLMs on resource-constrained devices presents unique challenges that differ significantly from traditional cloud-based implementations. These challenges span multiple areas, including model compression, efficient inference, security, energy efficiency, and seamless integration with diverse hardware platforms. Moreover, the dynamic nature of edge environments and the need for continuous adaptation introduce additional complexities that must be considered. We outline the most pressing challenges and opportunities in advancing the field of LLMs on-device. By identifying these key areas and stimulating innovation in developing more capable, efficient, and reliable on-device language models, we aim to provide insights for future research efforts. We should notice that the challenges and opportunities discussed here are interconnected: the progress in one area often has implications for others. Therefore, a holistic approach that considers the interplay between different aspects of on-device LLM deployment is essential for achieving significant advancements in the field. We delve into the current state of research, identifying key challenges and proposing potential directions for future work, summarized in Fig. 6. By addressing these challenges, researchers and practitioners can push the boundaries of what is possible with on-device LLMs, ultimately leading to more intelligent, efficient, and user-centric computing experiences across various applications and domains.